### v3

# Ultracoherent superconducting cavity-based multiqudit platform with error-resilient control

Abstract
Superconducting radio-frequency (SRF) cavities offer a promising platform for quantum computing due to their long coherence times, yet integrating nonlinear elements like transmons for control often introduces additional loss. We report a multimode quantum system based on a 2-cell elliptical shaped SRF cavity, comprising two cavity modes weakly coupled to an ancillary transmon circuit, designed to preserve coherence while enabling efficient control of the cavity modes. We mitigate the detrimental effects of the transmon decoherence through careful design optimization that reduces transmon-cavity couplings and participation in the dielectric substrate and lossy interfaces, to achieve single-photon lifetimes of 20.6 ms and 15.6 ms for the two modes, and a pure dephasing time exceeding 40 ms. This marks an order-of-magnitude improvement over prior 3D multimode memories. Leveraging sideband interactions and novel error-resilient protocols, including measurement-based correction and post-selection, we achieve high-fidelity control over quantum states. This enables the preparation of Fock states up to $N=20$ with fidelities exceeding 95%, the highest reported to date to the authors’ knowledge, as well as two-mode entanglement with an estimated coherence-limited fidelities of 99.9% after post-selection. These results establish our platform as a robust foundation for quantum information processing, allowing for future extensions to high-dimensional qudit encodings.
I introduction
Bosonic quantum systems have emerged as a compelling platform for quantum computing, offering several intrinsic advantages over traditional qubit-based architectures. By encoding quantum information in harmonic oscillator modes, these systems benefit from long coherence times, large Hilbert spaces, and relatively simple error channels dominated by photon loss. Coupling to nonlinear superconducting circuits has enabled universal control and gate operations using a variety of techniques [SNAP2015PRL, Eickbusch2022, rosenblum2018cnot, heeres2017implementing, gao2019entanglement, PRXQuantum.4.020355, LuSchoelkopf2023, Maiti2025] in both single-mode [reagor2016quantum] and multimode [wang2016schrodinger, Chakram2021seamless] circuit QED systems implemented with three-dimensional (3D) SRF cavities featuring millisecond-scale coherence times. These capabilities have enabled hardware-efficient implementations of bosonic quantum error correction, including demonstrations that surpass the break-even point [ofek2016extending, sivak2023real, ni2023beating]. More recently, these control strategies have begun to extend to ultra-high-coherence cavities [Rosenblum2023mushroom, Oriani2024Nb_coax], with coherence times exceeding tens of milliseconds and the potential for further improvements [Romanenko2020], opening new avenues for scalable quantum information processing.
Qubit-based quantum computing has long been the dominant paradigm. In bosonic error correction, cavity modes are typically used to encode logical qubits [mirrahimi2014dynamically, michael2016new]. The large Hilbert space of a cavity also makes it a natural platform for implementing a qudit using multiple quantum levels, both natively and with error correction [brock2025quantum]. Qudits offer enhanced information density and can enable more compact and efficient quantum circuits across a wide range of algorithms [bullock2005asymptotically, wang2020qudits, low2025]. They are particularly well-suited for quantum simulations in chemistry [cao2019quantum, dutta2024simulating], condensed matter [Chizzini2024], and high-energy physics (HEP) [Daniel2022, Zache:2023cfj], where they can represent physical degrees of freedom such as large spins, rotors, and gauge fields [Girvin2024hybrid]. In HEP applications, a quantum field can also be discretized in position space and encoded within a single high-dimensional qudit [gustafson2022primitive]. The emergence of high-coherence cavities presents a promising path for improved control over high-photon-number Fock states, thereby enabling qudit-based quantum computing and simulation.
Despite their advantages, high-coherence cavities pose increasing challenges for quantum control as coherence times improve. A central challenge lies in balancing fast, high-fidelity control and preserving the cavity’s exceptional coherence. Strong coupling to ancillary transmon circuits enables fast gates but introduces photon loss through the inverse Purcell effect and exposes the cavity to backaction from ancilla errors. One approach that resolves this tension is to reduce the coupling between the cavity and the lossy ancillary elements to protect cavity coherence, while overcoming the gate-speed limitations of weak coupling through techniques such as conditional displacements [Eickbusch2022], sideband interactions [huang2025fast], and sideband-enhanced selective number-dependent arbitrary phase (SNAP) gates [you2025floquet]. Nevertheless, as coherence continues to improve, it becomes increasingly important to develop control strategies that are not only fast but also fault-tolerant to the dominant errors in the system [pietikainen2024strategies].
In this work, we present a two-mode bosonic quantum system implemented within a multimode SRF cavity, achieving ultrahigh coherence times and novel cavity control capabilities resilient to ancilla decoherence. Our system utilizes a two-cell niobium SRF cavity design based on the TESLA geometry [Padamsee1998book, Aune2000], weakly coupled to a transmon circuit for precise quantum control while preserving cavity coherence. We achieve single-photon lifetimes of 20.6 $\pm$ 0.4 ms and 15.6 $\pm$ 0.2 ms for the two bosonic modes, significantly longer than most state-of-the-art cavity QED platforms, establishing a record for multimode quantum memories. Despite the coherence-preserving weak coupling, we employ sideband operations to achieve fast cavity control on timescales far shorter than those set by the weak dispersive shift. In conjunction with novel protocols that make cavity control leading-order immune to ancilla errors, we achieve high-fidelity preparation of Fock states up to $N=20$ with fidelity exceeding 95%, as well as two-mode entanglement with estimated coherence-limited fidelities up to 99.9% after post-selection. These advancements in coherence and control mark a significant step in bosonic quantum computing, laying the groundwork for exploring high-dimensional qudit encodings and gate operations.
II Cavity Characterization and System Design
We designed a two-mode superconducting cavity based on the multi-cell TESLA geometry [Aune2000, Reineri2023], consisting of two elliptically shaped cells connected by a circular aperture referred to as the iris, as shown in Fig. 1(a). The hybridization of the nearly degenerate bare TM010 modes in the two cells gives rise to two normal modes, named “Alice” and “Bob”, which have symmetric and antisymmetric electric field profiles, respectively (Fig. 1(b)). This normal mode structure leads to similar electric field density at the opening of the transmon tunnel, allowing for even coupling strength between the transmon and the two modes. The radius of the iris controls the coupling strength and consequently the frequency detuning between the normal modes. The Alice and Bob modes are designed to have frequencies of 5.779 GHz and 6.872 GHz, respectively. This frequency separation helps minimize crosstalk during control operations between the cavity modes, while also providing flexibility in placing the transmon frequency, which can vary due to fabrication inaccuracies. Cylindrical waveguide pipes with sufficiently high cut-off frequencies (35 GHz) are attached on either side of the cavity; one of these houses the transmon. Additionally, each pipe includes two ports for microwave drive lines to control transmon, readout and cavity modes.
Figure 1: Device architecture and coherence properties. (a) Schematic of the double-cell elliptical SRF cavity with four ports for RF control. Three ports are used in this experiment. The chip (blue) contains an ancilla transmon, a Purcell filter, and a stripline resonator for measuring the transmon. (b) Picture of the niobium cavity (top panel) with mounting brackets made of oxygen-free copper. Bottom panel shows the electric field distribution of the two fundamental modes dubbed ‘Alice’ and ‘Bob’. (c) Coherence times of the single photon states obtained using T1 and Ramsey experiments. (d) Relaxation times of Alice mode’s Fock levels up to $|20\rangle$ with a fit to $1/(n\gamma_{1})$, where $n$ is the Fock state photon number and $\gamma_{1}=1/T_{1}^{(1)}$ is the decay rate of $|1\rangle$. The inset shows that the theoretical decay rate of Fock $|n\rangle$ is equal to $n\gamma_{1}$ for a harmonic oscillator.
We manufactured this cavity from high-purity niobium, using e-beam welding to eliminate seam loss. State-of-the-art surface treatments, such as chemical etching and heat treatments, were applied to remove damaged surface layers and contaminants. The optimized surface quality and the intrinsic bulk property of high-purity niobium enable high quality factors for the cavity modes in the milli-Kelvin and low-photon-number regime [Romanenko2020, Paul2021, Rosenblum2023mushroom, Oriani2024Nb_coax, charpentier2025], providing an excellent platform for quantum computing applications.
To establish a baseline for the cavity’s performance before integrating the transmon qubit, we first conduct a milli-Kelvin characterization of the bare cavity in a dilution refrigerator. Using a classical decay measurement, we confirm that the two fundamental modes exhibit exceptionally long decay times of 25.5 ms and 19.3 ms, respectively, corresponding to loaded quality factors of approximately $9.4\times 10^{8}$ and $8.4\times 10^{8}$. We keep the probe antenna highly undercoupled ($Q_{\rm ext}>10^{11}$), thus preventing radiation losses from limiting the cavity lifetime. We also perform temperature-dependent measurements of the cavity’s decay times to distinguish contributions from residual resistance and two-level systems (TLS), obtaining results comparable to established SRF best practices (see Appendix B). These findings confirm the suitability of this design for quantum memory applications and demonstrate that our fabrication and preparation steps effectively preserve the cavity’s coherence, laying the foundation for subsequent transmon integration.
To achieve universal control of the two-mode cavity as a quantum processing unit (QPU), we introduce a transmon qubit that simultaneously couples to the two normal modes, thanks to their similar electric field strengths at the opening of the transmon tunnel. This integration requires careful optimization since the transmon can introduce inverse Purcell decay, dielectric loss and seamloss at the cavity-chip holder interface. We adopt two main strategies for this optimization. First, we carefully engineer the cavity geometry to minimize the transmon chip’s participation in the cavity field. Second, we reduce the dispersive shift by adjusting the transmon’s location along the axis of the cavity, further lowering its participation ratio while maintaining sufficient coupling for cavity control (see Appendix C). RF simulations confirm that these modifications preserve the TESLA geometry’s inherent low-loss advantage, enabling robust two-mode control via the ancilla qubit without sacrificing the cavity’s long coherence times.
[…]
A characteristic property of a linear system is that the decay rate of a Fock state $\gamma_{n}$ is proportional to its photon number, i.e., $\gamma_{n}=n\gamma_{1}$. Equivalently, the lifetime of the $n$-th Fock state satisfies $T_{1}^{(n)}=T_{1}^{(1)}/n$. We experimentally verify this relationship by preparing Fock states with increasing photon numbers (as described in the next section) and measuring their relaxation times. Figure 1(d) shows the measured data in excellent agreement with the theoretical expectation. Remarkably, the lifetime of the $|20\rangle$ state remains above 1 ms, demonstrating the system’s potential as a high-coherence platform supporting Fock states with large photon numbers, a precursor to qudit control.
III Ancilla-Error-Resilient Cavity State Preparation
An ancilla qubit dispersively coupled to a linear system enables universal control [SNAP2015PRA, law1996arbitrary, Girvin2024hybrid]. The resulting photon-number dependence of the ancilla transition frequency enables so-called SNAP gates, in which arbitrary phases can be applied to specific Fock states. Combined with cavity displacements, this allows for universal control of the cavity state [SNAP2015PRA, SNAP2015PRL]. However, the standard SNAP gate duration should be sufficiently longer than $2\pi/\lvert\chi_{e}\rvert$, making it inefficient when $\chi_{e}$ is small. In contrast, both the Echoed Conditional Displacement (ECD) gate [ECD2022NatPhys, Diringer2024CNOD, You2024crosstalk] and sideband control [sideband2015PRA, Liu2021qudit, huang2025fast] also enable high-fidelity universal control, with a gate time controlled by microwave drive strength much shorter than the SNAP gate duration limit $2\pi/\lvert\chi_{e}\rvert$.
In this experiment, we utilize the sideband (SB) scheme [sideband2015PRA, law1996arbitrary, huang2025fast] for controlling the cavity and further develop ancilla-error resilient state preparation schemes. The basic principle of preparing a large Fock state in a single mode is illustrated in Fig. 2(a). The transmon-resonator system is initially in state $|g,0\rangle$. The transmon is unconditionally excited to the $|f\rangle$ level applying broadband $\pi_{ge}$ and $\pi_{ef}$ pulses. Then a sideband $\pi_{\mathrm{sb}}^{(0)}$ pulse at the $|f,0\rangle\leftrightarrow|g,1\rangle$ transition is applied to prepare the state $|g1\rangle$ (see Appendix D for the calibration of the sideband pulse). By repeatedly applying a sequence of $\pi_{ge}$, $\pi_{ef}$, and $\pi_{\mathrm{sb}}^{(n)}$ sideband pulses at $|f,n\rangle\leftrightarrow|g,n+1\rangle$, it is possible to prepare Fock states up to a large photon number. It’s worth noting that the sideband pulses need not be frequency selective for Fock state preparation, or even for certain superposition states utilizing a recently developed shelving technique [huang2025fast]. The sideband operation also allows for driven-dissipative reset of the cavity state in a time scale much faster than its natural lifetime (see Appendix E), which is utilized throughout our experiment.
In order to estimate the state preparation fidelity $\mathcal{F}_{n}$ of a Fock state with photon number $n$, one can avoid full tomography as $\mathcal{F}_{n}$ is equal to the population of $|n\rangle$. We measure the cavity Fock state occupation using photon-number-resolved spectroscopy (PNRS), where the probabilities are normalized using the transmon’s readout ($|g\rangle$ and $|e\rangle$) values. The population distribution for the Fock state 20 is shown in the first panel of Fig. 2(c), with peak probability around 0.51 limited by the leakage to the lower Fock levels, as a result of decoherence and control errors occurring during the SB protocol.
In climbing the Fock ladder to reach the target photon state, decoherence in both the transmon and cavity can introduce errors during state preparation. Because the cavity coherence time significantly exceeds that of the transmon, and since transmon $\pi$ pulses in our system are approximately twenty times faster than sideband operations, the fidelity is primarily limited by transmon errors occurring during sideband pulses. Specifically, a transmon decay event from the initial state $|f,n\rangle$ results in the error state $|e,n\rangle$, preventing the sideband transition to $|g,n+1\rangle$; similarly, transmon dephasing disrupts the coherent transition $|f,n\rangle$ to $|g,n+1\rangle$, leaving residual population in $|f,n\rangle$. Although the probability of such events within a single sideband operation is low, the multiple repetitions required to prepare high-photon-number Fock states can cumulatively lead to significant infidelity. To mitigate this cumulative infidelity caused by the ancilla, we have developed a “sideband feedforward protocol” (SFP), illustrated in Fig. 2(b). Akin to other techniques exploiting multi-level transmon readout for fault-tolerant cavity operations [Rosenblum2018, Elder2020, Reinhold2020, Taka2023, teoh2023dual], the SFP leverages error syndrome detection via high-fidelity single-shot measurements of the transmon populations in the $|g\rangle$, $|e\rangle$, and $|f\rangle$ states after each sideband operation. This information is then fed forward to determine the subsequent corrective pulse in real time. A measurement outcome of $|f\rangle$ heralds a dephasing event, which we correct by repeating the same sideband pulse. An outcome of $|e\rangle$ indicates a transmon decay event; in this case, we first apply a $\pi_{ef}$ pulse and then repeat the sideband pulse to restore the intended state. Finally, an outcome of $|g\rangle$ indicates the ideal scenario where no error has occurred, and thus requires no further correction.
Figure 2: Fock state preparation and characterization. (a) Sideband scheme demonstrating the transfer of population starting from $|g0\rangle$ to $|gn\rangle$. Addition of one photon in the bosonic ladder involves application of unconditional $\pi_{ge}$ (light blue arrows) and $\pi_{ef}$ (red arrows) pulses on the transmon followed by a $\pi$ pulse ($\pi_{sb}$) which allows the sideband transition $|f,n\rangle\leftrightarrow|g,n+1\rangle$ (green arrows). (b) Pulse protocol for correcting ancilla errors and cavity single-photon loss. After each $\pi_{sb}$ pulse, the transmon is measured, followed by application of conditional pulses to correct transmon errors, termed as ‘sideband feedforward protocol’ (SFP). At the end of the sequence, a parity filter (PF) is applied to post-select results with the expected parity. The PF is implemented by applying two opposite $\pi_{ge}/2$ pulses conditioned on $n$ photons with a gap of $\pi/|\chi_{e}|$ so that the correct state is always mapped to $|g\rangle$. (c) Photon-number-resolved spectroscopy (PNRS) of the $\lvert 20\rangle$ state, prepared in the Alice mode, using sideband (SB) only (left), SFP (middle), and both SFP and PF (right). SFP significantly improves the height of the target peak, with remaining infidelity primarily caused by leakage to the off-target state $\lvert 19\rangle$. This error is further suppressed by implementing PF and post-selecting on the correct state. (d) Fidelities of different Fock states in the Alice (round markers) and Bob (triangular markers for $|20\rangle$ state only) modes prepared using the three methods. The error-resilient methods clearly improve the state preparation fidelities.
[…]
Furthermore, we observe that the remaining off-target population after climbing the ladder to Fock state $N$ using SFP is mostly found in $|N-1\rangle$, making it possible to apply a parity filter (PF) and post-select out the $|N-1\rangle$ population for an additional improvement. Specifically, we apply a $\pi_{ge}/2$ pulse at the frequency $\omega_{0}+N\chi$, wait for $\pi/|\chi_{e}|$, then apply a $(-\pi_{ge})/2$ pulse, and finally perform a transmon readout. During the wait time of $\pi/|\chi_{e}|$, the transmon acquires a phase shift dependent on whether the total photon number in the cavity is even or odd. The final $(-\pi_{ge})/2$ pulse maps correct parity states to $|g\rangle$ and incorrect parity states to $|e\rangle$. We then discard measurement outcomes where the transmon is found in $|e\rangle$, indicating a parity error. By combining the PF with SFP, we further suppress the off-target population, resulting in a higher preparation fidelity (Fig. 2(c) right panel).
We perform an open-system simulation for the Fock state preparation protocol, and details are provided in Appendix H. The simulation shows that SFP effectively corrects the transmon decoherence errors, which are the main source of infidelities in the SB case. While the SFP fidelity is noticeably improved by correcting the transmon decoherence errors, the remaining infidelity arises primarily from transmon readout errors and qudit relaxation errors caused by the additional readout operations. This change in the source of infidelities explains the out-of-target population predominantly observed in the state $|N-1\rangle$ when using SFP compared to a more uniform out-of-target distribution in the SB case, as shown in Fig. 2(c), since the effective qudit relaxation rate increases with the photon number while the transmon decoherence errors are largely insensitive to the photon number. The simulation results indicate that the fidelities of SFP can be further enhanced by optimizing transmon readout and reducing qudit decoherence, particularly when preparing high-photon-number Fock states with large effective qudit relaxation rates.
We plot the state preparation fidelities of Fock states up to twenty photons for Alice and Bob modes in Fig. 2(d). The traces show clear improvement of the Fock state fidelity obtained with SFP and PF techniques, compared to the vanilla SB protocol. We achieve preparation fidelities of 95.3±1.9%95.3\ $\pm$ 1.9\ $\%$ and 94.6± 1.1%94.6\ $\pm\$$1\%$ for the Fock state $|\psi\rangle_{A/B}=|20\rangle$ in the Alice and Bob modes, respectively. To the best of our knowledge, this represents the highest reported fidelities for such high-photon-number Fock states to date. The high-fidelity Fock state preparation also enables calibration of sideband control at even higher photon numbers, preserving high contrast in the calibration measurement. While we have demonstrated only basis state preparation, sideband control enables arbitrary state synthesis [law1996arbitrary, hofheinz2009synthesizing] and theoretically supports universal multi-qudit gate operations [mischuck2013qudit, Liu2021qudit]. Achieving high fidelity and long coherence for large Fock states thus opens new opportunities for qudit-based quantum computation [wang2020qudits, Blok2021, Goss2022, Roy2023two_qutrit, nguyen2024empowering, Nikolaeva:2024wxl, Gustafson:2025buk, Litteken2023quantum_waltz, Cao2024transmon4, Wang2025transmon12, Blok2025transmon8, shi2025Grover_qudit], quantum sensing [Agrawal2024stimulated, deng2024metrology], and quantum simulation [gustafson2021prospects, ciavarella2021trailhead, Zache:2023cfj, bauer2023quantum, gustafson2022primitive, illa2023quantum, Kurkcuoglu:2024cfv, Girvin2024hybrid, meth2025LG_qudit].
IV Sideband-mediated Virtual Raman Interaction for Cavity Entanglement
Figure 3: Virtual Raman-assisted beamsplitter (VRBS) operation. (a) Level diagram showing microwave drives for realizing the VRBS operation between the two modes and relevant stochastic processes that can lead to errors. The green arrows represent detuned $|f0\rangle\leftrightarrow|g1\rangle$ transition for both modes with rates $g_{sb}$. The gray lines show single-photon decay channels giving rise to erasure errors. The red lines represent transmon heating events arising due to sideband drives. (b) Characterization of VRBS interactions. The three panels show beamsplitter rates, cavity decoherence rates, and gate fidelities as a function of the sideband detuning. Blue and red markers denote experimental data and numerical simulation results, respectively. (c) The time-domain VRBS measurement at the optimal detuning. The top panel shows the circuit diagram, while the middle and bottom panels show data before and after discarding shots containing the erasure error. Fidelity numbers are obtained using Eq. (2). (d) Circuit diagram and the three oscillation graphs showing the VRBS-induced swap operations with heating checks. The swap operations $U_{\text{SWAP}}$ are applied to the initial state, with a swap time determined from the result in (c). The transmon is measured after each swap (top oscillation graph) and a heating check is performed by not discarding the shots only if it is found in $|g\rangle$. The gates $U_{\rm BS}$ and $U_{\rm BS}(\phi)$ are only applied for the initialization and readout of the state $(|10\rangle+|01\rangle)/\sqrt{2}$. The extracted swap fidelity improves when the heating check is performed, for both the $|10\rangle$ and $(|10\rangle+|01\rangle)/\sqrt{2}$ initial states (middle and bottom graph).
As a demonstration of two-mode control in this cavity system, we extend the sideband approach to implement a two-mode entangling operation, realizing a beamsplitter interaction in the single-photon subspace of the two modes. Such an interaction has previously been implemented using either a transmon [Gao2018, pfaff2017controlled] or a dedicated coupling element to enable three- or four-wave mixing between modes [PRXQuantum.4.020355, LuSchoelkopf2023]. In contrast to standard four-wave mixing with a fixed-frequency transmon, we leverage the transmon’s $|f\rangle$ level to mediate the interaction. By simultaneously driving the transmon at a detuning $\Delta$ from the $|f0\rangle\leftrightarrow|g1\rangle$ sideband transitions of both Alice and Bob, we activate a virtual-Raman-assisted beamsplitter (VRBS) interaction between the two modes, enabling coherent photon exchange and entanglement (see Fig. 3(a)). Choosing $\Delta$ larger than the sideband rates suppresses direct population of the more lossy transmon $|f\rangle$ state during the interaction, albeit at the cost of a reduced conversion rate. Importantly, compared to the standard four-wave mixing protocol, our technique achieves a faster conversion rate for a given drive strength, with an enhancement factor of $|\alpha/\Delta|$, where $\alpha$ is the transmon anharmonicity (see Appendix I).
To estimate the fidelity of the entangling operation, we measure the evolution of the two modes under the VRBS interaction, with an initial single-photon Fock state in the Alice mode. After an evolution time of $t$, we turn off the VRBS interaction, and map the two-mode states in the single-photon subspace, specifically $\rangle, |01|01\rangle$, and the erasure state $|00\rangle$ (order: Alice $\otimes$ Bob), onto the transmon states $|e\rangle$, $|f\rangle$, and $|g\rangle$, respectively (see Appendix M). This allows us to monitor the evolution of the state populations as a function of time, which can be fitted to the theoretical model described in [LuSchoelkopf2023],
$P_{\text{Bob}}=\frac{1}{2}e^{-\kappa_{1}t}(1+e^{-\kappa_{\phi}t}\cos(2g_{\mathrm{BS}}t)),$ (1)
where $g_{\mathrm{BS}}$ is the beamsplitter rate, with the corresponding beamsplitter time $\pi/(4g_{\mathrm{BS}})$. The effective decay and dephasing rate of the oscillations, $\kappa_{1}$ and $\kappa_{\phi}$, together determine the upper-bound of fidelities ($F^{*}_{\mathrm{BS}}$, $F^{*}_{\mathrm{SWAP}}$) of the beamsplitter and the swap operations. We refer to them as coherence-limited fidelities, which are given by
$F^{*}_{\mathrm{BS}}\approx 1-\frac{\pi}{4}\frac{\kappa_{\mathrm{BS}}}{g_{\mathrm{BS}}},\;F^{*}_{\mathrm{SWAP}}\approx 1-\frac{\pi}{2}\frac{\kappa_{\mathrm{BS}}}{g_{\mathrm{BS}}},\;\kappa_{\mathrm{BS}}=\kappa_{1}+\frac{\kappa_{\phi}}{2}.$ (2)
We optimize these fidelities as a function of the VRBS detuning $\Delta$, under fixed sideband drive strengths, $g_{sb}^{a}/2\pi=269.5\pm 0.5$ kHz, $g_{sb}^{b}/2\pi=301.8\pm 0.5$ kHz. The Raman-assisted beamsplitter rate increases as $\Delta$ decreases, as shown by the experimental data in Fig. 3(b, left). However, this enhanced interaction comes at the cost of both increased decay and dephasing. Qualitatively, the beamsplitter rate scales inversely with $\Delta$ (see Appendix I), while the sideband-drive-induced Purcell decay follows an approximate $\Delta^{-2}$ dependence [you2025floquet]. This scaling alone would suggest that the total infidelity should decrease with increasing detuning. However, at large detunings, cavity decoherence is no longer limited by sideband-induced decay, but is instead dominated by dephasing due to transmon heating induced by the strong drive tones. This heating arises from undesired parametric processes such as dressed dephasing [Boissonneault2008]. A heating event excites the transmon to the $\lvert e\rangle$ state, effectively halting the coherent photon exchange. These random transmon heating events dephase the VRBS oscillation as a result of the strong transmon-state-dependence of the beamsplitter rate (see Appendix I). Additionally, transmon heating enhances photon shot noise dephasing of the cavity modes. Since these heating processes are governed by the detuning between the transmon and the drive (on the order of GHz) [yaxing], they are largely insensitive to MHz-scale variations in $\Delta$, leading to the observed saturation of the beamsplitter dephasing time $1/\kappa_{\phi}$ at large detunings in Fig. 3(b, middle).
The interplay between these mechanisms leads to a maximum in gate fidelity as a function of sideband detuning, as shown in Fig. 3(b, right). We quantitatively capture this trade-off using open-system numerical simulations, from which we extract cavity decoherence times, beamsplitter rates, and the resulting gate fidelities across a range of detunings. The simulated results show good agreement with experimental measurements (see Appendix J for further discussion).
The coherent photon exchange process corresponding to the VRBS interaction at the optimal parameters are shown in Fig. 3(c, top). Fitting this oscillation using Eq. (2) yields a coherence-limited beamsplitter fidelity of $99.686\pm 0.004\%$. This operation can be interpreted as a single-qubit gate on a dual-rail qubit ($\{|10\rangle,|01\rangle\}$) encoded across the single-photon subspace of the two cavity modes [teoh2023dual, chou2024superconducting]. We observe a gradual increase in the population of the $|00\rangle$ state during the interaction as a result of photon loss in the cavity modes, equivalent to the erasure error of a dual-rail qubit. While we do not implement an erasure syndrome measurement that preserves the dual-rail qubit state, we estimate the improvement in fidelity from erasure checks by post-selecting on the absence of erasures at the end of the oscillation. The resulting oscillation, conditioned on no erasure events, exhibits a prolonged decay constant of $1/\kappa_{1}=70.85$ ms, translating to an improved $\mathcal{F}^{*}_{\mathrm{BS}}$ of $99.810\pm 0.003\%$ (Fig. 3(c, bottom)).
In addition to post-selecting on erasure errors, we also perform mid-circuit detection of transmon heating errors, which, as previously discussed, are the dominant source of dephasing in the VRBS operation. These heating errors are detected via a transmon measurement, which—when post-selecting on the transmon being in the ground state—introduces negligible backaction on the cavity states due to the weak hybridization between the transmon and the cavities in our regime. We verify that the cavity Fock state lifetimes are unaffected by repeated transmon measurements in Appendix M. We demonstrate the resulting fidelity improvement using a SWAP operation. To amplify the effect of heating, we apply a sequence of SWAP gates implemented via the same VRBS interaction. As a baseline, Fig. 3(d, top) shows the outcome of this sequence—initialized in $\lvert g10\rangle$—with only a final erasure check. The measured heating rate is found to be approximately 1% per SWAP (See Appendix L). To quantify the benefit of heating error detection, we then perform transmon measurements after each SWAP. The corresponding results, also shown in Fig. 3(d, middle), exhibit a clear improvement in contrast after one hundred SWAPs, corresponding to a two-fold improvement in fidelity (infidelity drops from $0.4\%\rightarrow 0.2\%$). We further verify that this improvement holds for initial states prepared as entangled superpositions of the two modes (i.e., equatorial states on the Bloch sphere in the dual-rail encoding) by applying a $U_{\mathrm{BS}}$ following the initialization of a Fock state in one mode (Fig. 3(d, bottom)), and another $U_{\mathrm{BS}}(\phi)$ before the readout to account for Stark-shift-induced phase rotations (see Appendix K). We thereby realize a partially error-detected SWAP gate, which can be incorporated mid-circuit to improve the fidelity of shallow-depth algorithms and to support dual-rail surface code implementations [Taka2023] using simpler fixed-frequency transmon couplers. This error-detection scheme can also be extended to beamsplitter operations via additional pulse calibration steps that compensate for Stark shift effects [LuSchoelkopf2023]. Based on the SWAP gate measurement results, we infer a coherence-limited fidelity approaching $99.9\%$ for the error-detected beamsplitter gate—comparable to state-of-the-art results, albeit with longer gate durations, but achieved using significantly simpler circuit hardware.
The demonstrated VRBS operation entangles the two cavities within the single-photon subspace. When extended beyond this subspace, the interaction continues to couple states of fixed total photon number and mediates single-photon exchange between the modes. However, the nonlinearity of the transmon introduces photon-number-dependent corrections to the beamsplitter rate, as shown in Appendix I. In particular, the rate is modified because the VRBS detuning for higher-photon-number sectors is altered by their dispersive shift. This nonlinear (non-Gaussian) beamsplitter can serve as a primitive for implementing entangling gates between qudits and for quantum simulations of interacting bosonic systems. For example, in the two-qutrit subspace, five VRBS operations combined with single-qutrit rotations can be used to synthesize the qutrit CSUM gate with 99.1% fidelity (see Appendix N).
V Conclusion
We have demonstrated a superconducting quantum platform that combines ultrahigh coherence with error-resilient control in a two-mode SRF cavity-based circuit QED system. By leveraging sideband interactions and a real-time feedforward protocol, we achieve high-fidelity preparation of Fock states up to $n=20$, with post-selected fidelities exceeding 95%. We further realize two-mode entangling operations mediated by a virtual Raman transition through the intermediate $|f00\rangle$ state, allowing for coherence-limited beamsplitter fidelities up to 99.9% after post-selection. The current gate speed is limited by the sideband rate, which can be improved through further design and wiring optimizations. Extending beyond the single-photon manifold, the ancilla-mediated interaction introduces nonlinear terms in the effective Hamiltonian that may be harnessed for simulating interacting bosonic systems and as primitives for universal qudit control—directions we plan to explore in future work. We also aim to investigate the application of feedforward in the context of gate operations.
The elliptical-shaped 2-cell cavity architecture demonstrated in this work can be naturally scaled by increasing the number of cells to realize larger multimode systems. These can be configured into cascaded random access quantum memories [li2025cascaded] with order-of-magnitude improvements in coherence times, where each two-cell module serves as a cache memory. Improved surface treatments [Romanenko2020, Posen2020] can further improve quality factors by an order of magnitude. Additionally, these two-cell devices can function as modular building blocks for a scalable quantum module, leveraging quantum-state router [zhou2022modular] and state-of-the-art tunable coupling devices [LuSchoelkopf2023, PRXQuantum.4.020355, Maiti2025] for the exchange of quantum information and the distribution of entanglement. This setup can be further scaled by employing remote interconnection between the scalable modules, and even across different cryogenic systems, facilitating the development of a distributed quantum network with improved coherence and transformative computational power.
Acknowledgements.
The authors acknowledge Edward D. Pieszchala, Tim Ring, Davida Smith, Damon J. Bice, Michael H. Foley, Ryan Treece, Charles J. Grimm, Scott D. Adams, Theodore Ill, Mackenzie Ring, and Dominic Baumgart for their support in cavity fabrication, processing, and assembly; David I. Schuster for providing the JPA for the measurement; Kevin A. Villegas Rosales, Michal Goldenshtein, Taekwan Yoon, and Vikrant Mahajan for their valuable assistance with instrument setup and coding for measurements; and Zachary Goff-Eldredge and John W. O. Garmon for insightful discussions. The work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems (SQMS) Center under the contract No. DE-AC02-07CH11359. This work made use of the Pritzker Nanofabrication Facility, part of the Pritzker School of Molecular Engineering at the University of Chicago, which receives support from Soft and Hybrid Nanotechnology Experimental (SHyNE) Resource (NSF ECCS-2025633), a node of the National Science Foundation’s National Nanotechnology Coordinated Infrastructure [RRID: SCR_022955].
Appendix A System Hamiltonian parameters
In this experiment, the relevant bosonic modes include two 3D high-Q cavity modes (Alice and Bob), one quasi-planar readout resonator mode, and the weakly nonlinear ancilla transmon mode. The system Hamiltonian can be expressed as
[…]
Values of the system parameters are shown in Table A1. The inverse-Purcell limit on the cavity mode lifetime is determined by both the relaxation as well as the dephasing channel of the transmon, which can be approximated as [Rosenblum2023mushroom]
$T^{\mathrm{Purcell}}_{1a,b}\approx\left(\frac{\Delta_{a,b}}{g_{a,b}}\right)^{2}\frac{T_{2E}^{ge}}{2}.$ (A.2)
From this, we estimate decay times of $T^{\mathrm{Purcell}}_{1a}=1.17\,$s and $T^{\mathrm{Purcell}}_{1b}=0.34\,$s, far exceeding the bare Alice and Bob lifetime limited by the cavity’s intrinsic material loss. It’s worth noting that this approximation assumes the same dephasing rate measured from the Hahn-echo experiment for the actual dressed-dephasing-induced cavity decay process. In reality, such a decay process is sensitive to dephasing noises at the detuning between the transmon and the cavity, which can have very different spectral components than in the Hanh-echo experiment. A more precise understanding of the inverse-Purcell limit due to the dephasing noise would require probing the dephasing spectral density via spin-locking techniques [Yan2013].
We also estimate the cavity dephasing time limit due to the transmon thermal shot noise. In the small thermal population limit, we approximate the thermal-noise-induced cavity dephasing time as [Clerk2007]
$T^{\mathrm{th}}_{\phi a,b}=\frac{(\chi_{e}^{a,b})^{2}+(T_{1}^{ge})^{-2}}{\bar{n}^{q}_{\mathrm{th}}(T_{1}^{ge})^{-1}(\chi_{e}^{a,b})^{2}}.$ (A.3)
Plugging in parameters in Table A1, we obtain $T^{\mathrm{th}}_{\phi}\approx 59$ ms for both modes, reasonably close to the measured cavity dephasing times.
Parameter Description Value $\omega_{a}$ Alice mode frequency $2\pi\times 5.779\,\mathrm{GHz}$ $T_{1}^{A}$ Alice mode depolarization time $20.6\,\mathrm{ms}$ $T_{2}^{A}$ Alice mode Ramsey decoherence time $21.1\,\mathrm{ms}$ $T_{\phi}^{A}$ Alice mode pure dephasing time $43.2\,\mathrm{ms}$ $\bar{n}_{\text{th}}^{a}$ Alice mode thermal population $0.1\%$ $\omega_{b}$ Bob mode frequency $2\pi\times 6.872\,\mathrm{GHz}$ $T_{1}^{B}$ Bob mode depolarization time $15.6\,\mathrm{ms}$ $T_{2}^{B}$ Bob mode Ramsey decoherence time $20.5\,\mathrm{ms}$ $T_{\phi}^{B}$ Bob mode pure dephasing time $59.8\,\mathrm{ms}$ $\bar{n}_{\text{th}}^{b}$ Bob mode thermal population $0.1\%$ $\omega_{q}$ Transmon $|g\rangle\leftrightarrow|e\rangle$ frequency $2\pi\times 6.402\,\mathrm{GHz}$ $\alpha$ Transmon anharmonicity $2\pi\times 245\,\mathrm{MHz}$ $T_{1}^{ge}$ Transmon depolarization time $147.4\,\mu\mathrm{s}$ $T_{2}^{ge}$ Transmon Ramsey dephasing time $47.3\,\mu\mathrm{s}$ $T_{2E}^{ge}$ Transmon Hahn-echo coherence time 205.8 $\mu\mathrm{s}$ $\bar{n}_{\text{th}}^{q}$ Transmon thermal population $0.25\%$ $T_{1}^{f}$ Transmon $|f\rangle$-state lifetime $80.1\,\mu\mathrm{s}$ $T_{2}^{gf}$ Transmon $(|g\rangle+|f\rangle)$ coherence time $45\,\mu\mathrm{s}$ $g_{a}$ Transmon–Alice vacuum Rabi coupling $2\pi\times 5.841\,\mathrm{MHz}$ $g_{b}$ Transmon–Bob vacuum Rabi coupling $2\pi\times 8.114\,\mathrm{MHz}$ $\chi_{e}^{a}$ Transmon–Alice dispersive shift $2\pi\times 71\,\mathrm{kHz}$ $\chi_{e}^{b}$ Transmon–Bob dispersive shift $2\pi\times 96\,\mathrm{kHz}$ $\omega_{r}$ Readout resonator frequency $2\pi\times 8.379\,\mathrm{GHz}$ $T_{1}^{R}$ Readout resonator lifetime $266\,\mathrm{ns}$ $\chi_{e}^{r}$ Transmon–readout dispersive shift $2\pi\times 411\,\mathrm{kHz}$ $\omega_{p}$ Purcell resonator frequency $2\pi\times 8.529\,\mathrm{GHz}$ $E_{\text{J}}$ Josephson junction energy $2\pi\times 24.46\,\mathrm{GHz}$ $\chi_{r}^{a}$ Readout-Alice cross-Kerr $2\pi\times 34\,\mathrm{Hz}$ $\chi_{r}^{b}$ Readout-Bob cross-Kerr $2\pi\times 131\,\mathrm{Hz}$ $K_{a}$ Alice mode self-Kerr $2\pi\times 2\,\mathrm{Hz}$ $K_{b}$ Bob mode self-Kerr $2\pi\times 19\,\mathrm{Hz}$
Table A1: System parameters and their values measured or inferred from the experiments.
Appendix B Cavity fabrication and loss characterization
[…]
Figure A1: Bare cavity characterization. (a) Ringdown measurements for the Alice mode showing raw data and linear fits. (b) Similar measurements for the Bob mode. (c) Internal quality factors as a function of the mixing chamber plate’s temperature for both modes and fitting to the TLS model.
To assess the decay and dephasing time scales of the bare cavity modes in the milli-Kelvin regime, we perform ringdown measurements, applying a 200 ms pulse and recording the reflected signal with a vector network analyzer. As depicted in Fig. A1(a-b), we measure the emitted power from the cavity modes in two distinct ways–phase-insensitive (PI) and phase-sensitive (PS), using the methodology introduced in [Read2023]. The phase-insensitive trace is obtained by measuring the power amplitude (equivalent to the square sum of the cavity’s I and Q quadratures) at each time point, then averaging over many such measurements. The resultant decay trace is well described by an exponential decay function characterized by the cavity’s decay rate,
$P_{\text{PI}}\propto\langle I^{2}+Q^{2}\rangle$ $\propto e^{-\frac{t}{T_{1}}}.$ (B.1)
From fitting to the PI power traces, we obtain the energy relaxation times of $25.545\,\text{ms}\pm 9\,\mu\text{s}$ for the Alice mode and $19.299\,\text{ms}\pm 4\,\mu\text{s}$ for the Bob mode. The phase‑sensitive trace, on the other hand, requires the measurement of both amplitude and phase information of the cavity field. By ensemble averaging the $I$ and $Q$ quadratures separately, and then taking their squared sum, we obtain the PS trace that can be well-fitted by the following function,
$P_{\text{PS}}\propto\langle I\rangle^{2}+\langle Q\rangle^{2}$ $\propto e^{-\frac{t}{T_{1}}}\lvert\langle e^{i\theta(t)}\rangle\rvert^{2}\approx e^{-\frac{t}{T_{2}/2}}.$ (B.2)
Here, $\theta(t)$ is the random phase noise responsible for cavity dephasing. The last approximation in Eq. B.2 assumes frequency-independent spectral density for $\theta(t)$ over the relevant bandwidth. Under this assumption, the overall decay pf $P_{\text{PS}}$ reduces to a single exponential function, with a time constant of
$\frac{T_{2}}{2}=\left(T_{1}^{-1}+2T_{\phi}^{-1}\right)^{-1},$ (B.3)
where $T_{\phi}$ is the intrinsic dephasing time of the cavity. Therefore, the measurement of $T_{1}$ and $T_{2}$ times allows for the extraction of the pure dephasing times, which we find to be $(1700\pm 181)\,\text{ms}$ and $(807\pm 148)\,\text{ms}$ for the bare cavity modes. In these fittings, we only take the first 100 ms of the trace to minimize the influence of the noise floor.
Additionally, we perform a temperature sweep to investigate the temperature dependence of the internal quality factor ($Q_{0}(T)$). This allows for the extraction of the zero-temperature loss tangent ($\delta_{0}$) associated with two-level systems (TLS), the residual surface resistance ($R_{\text{res}}$), and the temperature measurement efficiency ($\beta$) using the following equation [Romanenko2020],
$\frac{1}{Q_{0}(T)}=F\delta_{0}\tanh\Big(\beta\frac{\hbar\omega}{2k_{B}T}\Big)+\frac{R_{\text{res}}}{G}.$ (B.4)
Here, $F$ is the surface oxide filling factor, and $G$ is the geometry factor,
$F=\frac{\int_{\text{ox}}\epsilon_{\text{ox}}|\vec{E}|^{2}dv}{\int_{\text{all}}\epsilon_{\text{0}}|\vec{E}|^{2}dv},$ (B.5)
$G=\omega\mu_{0}\frac{\int_{\text{all}}|\vec{H}|^{2}dv}{\int_{\text{sur}}|\vec{H}|^{2}d\sigma}.$ (B.6)
The Tesla cavity geometry offers a notably low surface oxide filling factor ($F<2\times 10^{-8}$) and a large geometry factor ($G\approx 270~\Omega$), beneficial for enhancing cavity photon lifetime. At each temperature point, we perform a ring-down measurement to extract the loaded quality factor, $Q_{L}$. We then calculate the internal quality factor $Q_{0}$, from this $Q_{L}$ in conjunction with the coupling quality factor $Q_{\text{ext}}$ obtained from circle fitting at base temperature. We then fit the $Q_{0}$ vs. temperature data to Eq. B.4 and extract $F\delta_{0}$, $\beta$, and $R_{\mathrm{res}}/G$ from the fit. The factors $F$ and $G$ are obtained from finite-element method simulations using Eq. B.5 and Eq. B.6, which allows us to determine $\delta_{0}$ and $R_{\mathrm{res}}$. The relevant parameters are shown in Table A2. The measured residual surface resistance is higher than that typically observed in standard SRF cavities [Posen2020, Romanenko2017, Romanenko2020], suggesting the presence of residual contaminants on the cavity surface. Subsequent studies with different cavities confirmed the likely presence of such contamination in the furnace used for the cavity preparation. Based on this, elimination of these additional losses can be achieved by an extra buffered chemical polishing (BCP) etching to remove the embedded contaminants. Furthermore, we plan to reduce the TLS losses coming from surface oxide via enhanced surface treatment techniques developed previously [Romanenko2020, Posen2020]. The most recent progress of SQMS (manuscript in preparation) includes demonstrating the ability to grow an air-exposure robust oxide with the drastically suppressed TLS density, which should allow a further multifold increase in the cavity quality factor. These combined improvements should allow at least an order of magnitude improvement in the energy relaxation times of both Alice and Bob modes.
$f_{0}$ (GHz) $G$ $(\Omega)$ $R_{\text{res}}(n\Omega)$ $\delta_{0} FF$ $\delta_{0}$
5.779 (Alice) 295 73.2 $8.0\times 10^{-10}$ $6.7\times 10^{-9}$ 0.12
6.872 (Bob) 298 116.5 $7.9\times 10^{-10}$ $1.5\times 10^{-8}$ 0.05
Table A2: Extracted parameters from fitting the TLS model.
Appendix C Analysis of transmon-induced cavity losses
To preserve the excellent coherence properties of the cavity modes, we carefully minimize transmon chip-induced losses when introducing the transmon as the ancillary control element. In our device, the transmon chip is inserted through a narrow tunnel at the end of the 2-cell cavity, a feature reminiscent of the “beampipe” found in accelerator-based TESLA cavities, where it serves as the passage for the particle beam. In our quantum device, however, this tunnel is engineered with a reduced diameter, forming a cylindrical waveguide with a high cutoff frequency well above the cavity’s fundamental modes. This design enables selective coupling between the cavity and the ancilla qubit, while effectively isolating the cavity field from other potential loss channels, such as radiation losses and seam losses that can arise with larger or open interfaces. To further support fast and high-fidelity qubit readout without compromising coherence, we incorporate an on-chip Purcell filter. This filter allows for a readout resonator linewidth of approximately $\kappa_{r}\approx 0.5$ MHz, while maintaining readout-limited linewidths of $\kappa_{q}\approx 20$ Hz for the qubit and $\kappa_{a,b}\approx 0.02$ Hz for the cavity modes.
Following the methodology outlined in [Ganjam2024], we use finite-element simulations to calculate the loss channel participation ratios $p_{\text{bulk}}$, $p_{\text{MA}}$, $p_{\text{SA}}$, and $p_{\text{MS}}$ for the cavity modes,
$p_{\text{bulk}}=\frac{\int_{\text{bulk}}\epsilon|\vec{E}|^{2}dv}{\int_{\text{all}}\epsilon|\vec{E}|^{2}dv},$ (C.1)
$p_{\text{MA}}=\frac{t_{\text{surf}}\int_{\text{MA}}\epsilon_{0}|\vec{E}|^{2}d\sigma}{\epsilon_{\text{r,MA}}\int_{\text{all}}\epsilon|\vec{E}|^{2}dv},$ (C.2)
$p_{\text{SA,MS}}=\frac{t_{\text{surf}}\int_{\text{SA,MS}}\epsilon|\vec{E}|^{2}d\sigma}{\int_{\text{all}}\epsilon|\vec{E}|^{2}dv}.$ (C.3)
Here, $\epsilon$ denotes the dielectric constant, $\epsilon_{0}$ is the vacuum dielectric constant, and $\epsilon_{r}$ is the relative dielectric constant, which we assume to be 10. Additionally, $t_{\mathrm{surf}}$ represents the thickness of the surface region, assumed to be 3 nm in this work.
Using these participation ratios and loss tangents, we estimate the cavity loss rate as $\kappa_{\text{bulk}}=\omega_{c}p_{\text{bulk}}\tan\delta_{\text{bulk}}$ from the bulk loss, and $\kappa_{\text{surf}}=\omega_{c}\sum_{i=\text{MS,MA,SA}}p_{i}\tan\delta_{i}$ from surface losses. We evaluate these loss rates as functions of the dispersive shift and transmon frequency, as summarized in Table A3. We target each dispersive shift for Alice and Bob to be within 50–100 kHz to minimize transmon chip-induced losses while maintaining sufficient coupling for fast sideband control. Our analysis indicates that, under the current parameter configuration, transmon chip-induced material losses do not significantly limit cavity $T_{1}$ performance.
Transmon Alice Bob $\omega_{ge}/2\pi$ (MHz) $\chi_{e}^{a}/2\pi$ (kHz) $\kappa_{\text{bulk}}/2\pi$ (Hz) $\kappa_{\text{surf}}/2\pi$ (Hz) $\chi_{e}^{B}/2\pi$ (kHz) $\kappa_{\text{bulk}}/2\pi$ (Hz) $\kappa_{\text{surf}}/2\pi$ (Hz) 6714 -21 0.017 0.004 -150 0.604 0.178 6434 -58 0.027 0.007 -68 0.101 0.014 6304 -120 0.036 0.009 -43 0.065 0.017
Table A3: Transmon and substrate induced loss rates for cavity modes.
The accurate determination of the transmon mode’s effect on the loss channel participation ratios of the cavity mode depends critically on properly accounting for the nonlinearity-induced corrections, especially when the Lamb shift is comparable to the transmon-cavity detuning. We employ the Energy Participation Ratio (EPR) approach [Minev2021] to obtain the Lamb-shifted transmon frequency, and then renormalize the linear inductance of the Josephson junction used in the RF participation ratio simulations accordingly. This way, we are able to capture the transmon-cavity detuning precisely, allowing for a reliable estimation of the cavity’s inherited loss from the transmon and the substrate material.
Appendix D Sideband pulse calibration
Figure A2: Sideband pulse calibration. (a) By initializing the state in $\lvert f,n\rangle$ and sweeping the sideband drive frequency, we observe the dip in the transmon $|f\rangle$ state population $P(f)$ that represents the corresponding sideband resonance frequency. (b) Driving the sideband transition at the measured resonance leads to the oscillations of $P(f)$, from which we extract the $\pi$-pulse time. We plot the first five sideband pulses for the Bob mode as an example, utilizing the SFP and PF techniques for enhancing the contrast of the calibration measurements.
The sideband drive in our experiment is defined as a flat-top pulse, with 40 ns squared sine-shaped ramp on each side to ensure adiabaticity. It is important to calibrate the resonance frequency as well as the $\pi$-pulse time for the sideband operation. Due to a drive amplitude-dependent Stark shift, the actual sideband transition frequency deviates from the bare resonance, $\omega_{c}-(2\omega_{q}+\alpha)$. To precisely determine the transition frequency, we sweep the drive frequency around the bare resonance, and measure the resulting population in the $|f\rangle$ state. The exact resonance is then identified by a noticeable drop in the $|f\rangle$ population (Fig. A2(a)). Once the precise transition frequency is identified, we measure time-domain Rabi oscillations between the $|f,n\rangle\leftrightarrow|g,n+1\rangle$ to calibrate the exact $\pi$-pulse duration (Fig. A2(b)). The calibration process can be combined with the SFP and PF protocols, which enables calibrating the high-photon-number sideband transition frequencies and pi pulses without losing contrast.
Appendix E Sideband-aided cavity reset operation
Figure A3: Efficient cavity reset operation using the sideband interaction. (a) The driven-dissipation process of the cavity population as a result of the cavity-transmon sideband interaction and the transmon decay. (b) Circuit diagram of the measurement for determining the duration of each sideband drive for the cavity reset operation. (c) Decay curves for $\alpha=4$ coherent state in the Alice mode, with (blue) and without (orange) using the sideband-aided reset protocol. The vertical line represents the duration of each sideband pulse for sufficiently resetting the cavity.
We realize an efficient cavity reset operation in our experiment that requires a much shorter time than via the cavity’s natural relaxation. As illustrated in Fig. A3(a), this reset operation is activated by a driven dissipation process, where the cavity population hybridizes with the transmon via the cavity-transmon sideband interaction, thus inheriting the much faster decay rate from the transmon. In our experimental regime where the sideband interaction strength is much greater than the transmon’s decay rate, the effective decay rate from $\lvert g,n\rangle$ to $\lvert e,n-1\rangle$ is around half of the transmon $\lvert f\rangle$ to $\lvert e\rangle$ decay rate [Lu2017]. Thus, the engineered driven dissipation, through the cascaded decay process of $\lvert g,n\rangle\rightarrow\lvert e,n-1\rangle\rightarrow\lvert g,n-1\rangle$, is happening at approximately the same rate as the transmon’s decay, orders of magnitude faster than the cavity’s relaxation. Using this protocol, we can reset an arbitrary initial cavity state to the vacuum state. Knowing the highest occupied Fock state $\lvert N_{\text{max}}\rangle$ in the cavity, we apply a sequence of sideband drives, going from $|g,N_{\text{max}}\rangle\leftrightarrow|f,N_{\text{max}}-1\rangle$ to $|g,1\rangle\leftrightarrow|f,0\rangle$, each lasting for a sufficiently long time comparing to the transmon’s decay time. Given its autonomous nature, this protocol does not require knowing the exact initial state, and can tolerate small detunings of the sideband drives from resonance.
Fig. A3(b, c) demonstrate the circuit diagram and the result of the sideband-assisted reset protocol for the $\alpha=4$ coherent state in the Alice mode. In this case, we initiate the sideband drive at $N_{\text{max}}=26$, as the cumulative population of all Fock states above $N_{\text{max}}$ is below 0.5%. We apply the sequence of sideband drives for different cavity photon numbers, sweeping the duration time of each pulse. We monitor the decay of the initial cavity state by measuring the cavity’s vacuum state population, via a photon-number-selective $\pi$-pulse conditioned at the zero-photon peak at the end of the whole sequence. Compared to the natural decay curve that reaches an average photon number below 0.005 within approximately 160 ms, the sideband-aided reset achieves the same photon number in only 25 ms, corresponding to a duration time of 1 ms for each sideband drive. This driven reset protocol can be further accelerated by parametrically coupling the cavity to the readout resonator with a large damping rate.
Figure A4: Three-level readout of the transmon. (a) IQ blobs and demarcation lines for $|g\rangle,|e\rangle$ and $|f\rangle$ states. (b) Resulting confusion (assignment) matrix.
Appendix F Transmon readout fidelity
The frequency, amplitude and length of the readout pulse is optimized to maximize the 3-state discrimination probabilities for the transmon. A Josephson parametric amplifier (JPA) with about 17.7 dB gain is used to amplify the readout signal at the base temperature. The IQ blobs and demarcation lines for the $|g\rangle,|e\rangle$ and $|f\rangle$ levels are pictured in Fig. A4(a). This mapping is used for making decisions in SFP for correcting ancilla errors during Fock state preparations. The corresponding confusion matrix is shown in Fig. A4(b) resulting in an overall readout fidelity of approximately 98.8%.
Appendix G Efficiency of the Sideband-Feedforward Protocol
Figure A5: Number of applied SFP operations for each Fock state. (a) Average number of applied sideband-feedforward protocol (SFP) operations during the preparation of Fock states in Alice’s mode. Ideally, the SFP is applied exactly once per sideband transition, assuming no errors occur (orange line). Experimentally, we find that the SFP is applied an average of $1.059$ times per desired sideband transition (blue line). (b) Measured average number of error syndrome states obtained during the SFP protocol for each targeted Fock state. (c) Measured success rate of the parity filter following the SFP operation. The filter rejects states with the incorrect parity, and the observed success rate is consistent with the Fock state preparation fidelity achieved using the SFP protocol alone.
The high coherence time of the transmon ancilla, relative to the sideband operation time, ensures low error probability per sideband operation, allowing for the dramatic improvement in the final state fidelity without drastically increasing the time cost using the SFP. We characterize the efficiency of the SFP by measuring the number of feedforward operations necessary for each step up the Fock state ladder (see Fig. A5). Experimentally, we apply SFP to prepare the cavity in the target Fock state, and measure how often additional feedforward operations are required. After repeating this measurement across many trials, we determine that feedforward is applied on average $1.059\pm 0.003$ times per ladder step, verifying minimal overhead. Additionally, we find that 0.49% of data shots cannot be corrected by the SFP, indicating errors outside the correctable subspace, such as photon decay events in the cavity. This low failure rate and minimal overhead confirm the high efficiency and robustness of our SFP implementation.
Appendix H Fidelity estimations for SFP
Here we present an open-system numerical simulation of the Fock state preparation protocol discussed in Sec. III. The simulation results indicate that the sideband feedforward protocol (SFP) nearly completely corrects decoherence errors in the ancilla transmon qubit. However, this comes at the cost of introducing additional errors due to transmon decoherence and measurement errors. This trade-off directly results in the dominant peak found in $|N-1\rangle$ of the out-of-target population, as shown in Fig. 2 and confirmed by simulation. In the following subsections, we will explain the noise model used in the simulation and discuss the insights the simulation results provide.
H.1 Noise model
In the following, we will discuss a noise model for the transmon readout error and the decoherence effects of both the qudit and the ancilla transmon qubit based on the experimental parameters presented in Table A1.
All unitary operations performed for the state preparation are rotations between two levels, $\hat{R}_{j,k}(\theta)$, where $j$ and $k$ are level indices of the qudit-transmon system and $\theta$ is the rotation angle. For example, the sideband transition is given as $\hat{R}_{(f,n),(g,n+1)}(\pi/2)$. Here, we construct the rotation unitaries as the qubit rotation around the $y$-axis between levels $j$ and $k$ while keeping other levels unchanged. This gives us
$\hat{R}_{j,k}(\theta)=$ $(|j\rangle\langle j|+|k\rangle\langle k|)\cos\frac{\theta}{2}+(|k\rangle\langle j|-|j\rangle\langle k|)\sin\frac{\theta}{2}$
$+\sum_{\ell\neq j,k}|\ell\rangle\langle\ell|.$ (H.1)
Alternative rotation unitaries can be constructed by, for example, taking the qubit rotation around the $x$-axis or assuming phase accumulation at other levels. Each alternative corresponds to a different pulse realization of the transition operation. However, the choice of construction does not affect the simulated fidelity of our Fock state preparation protocol. Therefore, we use the specific construction shown above for all rotations. Without noise, the rotation operations modify the system state $|\psi\rangle$ such that
$|\psi\rangle\rightarrow\hat{R}_{j,k}(\theta)|\psi\rangle.$ (H.2)
We use the Lindblad master equation to simulate the Fock state preparation protocol with noise. The Lindblad master equation, which governs the time evolution of the system density matrix $\rho(t)$, can be written as
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=-i[\hat{H}(t),\rho]+\sum_{\alpha}\mathcal{\hat{D}}[\hat{A}_{\alpha}]\rho.$ (H.3)
Here, $\mathcal{\hat{D}}[\hat{A}_{\alpha}]$ is the Lindblad superoperator given by $\mathcal{\hat{D}}[\hat{A}]\rho=\hat{A}\rho\hat{A}^{\dagger}-\tfrac{1}{2}\hat{A}^{\dagger}\hat{A}\rho-\tfrac{1}{2}\rho\hat{A}^{\dagger}\hat{A}$, each jump operator $\hat{A}_{\alpha}$ is associated with a decoherence channel to be discussed in the next paragraph, and $\hat{H}(t)$ is the system and control Hamiltonian generating the rotation operation such that
$\hat{R}_{j,k}(\theta)=\mathcal{T}\int^{T}_{0}\mathrm{d}t\ e^{-i\hat{H}(t)t},$ (H.4)
where $T$ is the gate time and $\mathcal{T}$ is the time-ordering operator. The noisy rotation operations $\mathcal{\hat{N}}_{j,k}(\theta)$ evolve the system density matrix such that
$\rho\rightarrow\mathcal{\hat{N}}_{j,k}(\theta)\rho.$ (H.5)
By assuming that the rotation angle is directly proportional to the gate time, $\mathcal{\hat{N}}_{j,k}(\theta)$ can be approximated with a second-order symmetric trotterization [Suzuki1991] such that
[…]
We determined numerically that a value of $m=4$ is adequate for our simulation.
Transmon depolarization and thermal heating for the ancilla transmon qubit with three levels are modeled by jump operators
$\hat{A}_{T_{1}^{q},\downarrow}=$ $\sqrt{\gamma_{T_{1}^{ge}}(1+\bar{n}_{\mathrm{th}}^{q})}(\sqrt{2}|e\rangle\langle f|+|g\rangle\langle e|),$ (H.8)
$\hat{A}_{T_{1}^{q},\uparrow}=$ $\sqrt{\gamma_{T_{1}^{ge}}\bar{n}_{\mathrm{th}}^{q}}(\sqrt{2}|f\rangle\langle e|+|e\rangle\langle g|),$ (H.9)
where $\gamma_{T_{1}^{ge}}=1/T_{1}^{ge}$ is the transmon decay rate, and $\bar{n}_{\mathrm{th}}^{q}$ is the average thermal population. For transmon pure dephasing, we use the three-level system model presented in [Li2012], in which pure dephasing is described by two jump operators
$\hat{A}_{T_{\varphi}^{q},f}$ $=\sqrt{2\gamma_{T_{\varphi}^{ge}}}|f\rangle\langle f|,$ (H.10)
$\hat{A}_{T_{\varphi}^{q},e}$ $=\sqrt{\gamma_{T_{\varphi}^{ge}}}|e\rangle\langle e|,$ (H.11)
where $\gamma_{T_{\varphi}}=1/T_{2}^{ge}-1/(2T_{1}^{ge})$ is the transmon pure dephasing rate. The factor $\sqrt{2}$ associated with the $f$ state comes as an approximation from the ratio between the noise sensitivity of the fgf-g energy difference and that of the ege-g energy difference, and more detailed discussions of the estimation of the dephasing time of the qubit can be found in [Koch2007, Groszkowski2018]. For the Alice-mode qudit, we write the jump operators for depolarization and pure dephasing as
$\hat{A}_{T_{1}^{A}}$ $=\sqrt{\gamma_{T_{1}^{A}}}\hat{a}$, (H.12)
$\hat{A}_{T_{\varphi}^{A}}$ $=\sqrt{\gamma_{T_{\varphi}^{A}}}\hat{a}^{\dagger}\hat{a}$, (H.13)
where $\gamma_{T_{1}^{A}}=1/T_{1}^{A}$ is the Alice-mode decay rate and $\gamma_{T_{\varphi}^{A}}=1/T_{2}^{A}-1/(2T_{1}^{A})$ is the Alice-mode pure dephasing rate. The jump operators associated with the Bob-mode qudit can be written similarly.
We also take into account the transmon readout fidelity, which introduces error to the SFP, as the correction is conditional on the ancilla transmon qubit readout. Although the confusion matrix shown in Fig. A4 is sufficient for readout error mitigation by inverting the Markov process modeling noise [Geller2020, Peters2023], it does not distinguish between transmon-relaxation-induced readout error, which SFP can partially correct, and classification error, which SFP cannot correct. To more accurately estimate the readout-related error, we model the readout fidelity using a transmon relaxation probability, denoted as $p_{\mathrm{readout}}$, along with symmetric classification error probabilities. We use this model to simulate the experiment that measures the confusion matrix, and then optimize both the transmon relaxation probability and the classification error probabilities such that the simulated results best match the experimentally obtained confusion matrix. This yields the following results: $p_{\mathrm{readout}}(e\rightarrow g)=0.0055$, $p_{\mathrm{readout}}(f\rightarrow e)=0.0110$, and the classification error rate array
$\begin{bmatrix}0.9976&0.0024&0.0000\\ 0.0024&0.9966&0.0010\\ 0.0000&0.0010&0.9990\end{bmatrix}$. (H.14)
Note that the experimentally obtained confusion matrix used for optimization is slightly different from the one shown in Fig. A4 due to parameter drift. In the simulation, we also take into account the qudit’s decoherence during the transmon readout by including the corresponding depolarization and dephasing effects over the readout time of approximately $1.7\,\mu\text{s}$.
H.2 Simulation results
Figure A6: Fock state preparation simulation for the Alice-mode qudit. (a) Simulated qudit population to prepare $|n=20\rangle$ Fock state. (b) Simulated fidelities improved by SPF and further improved by PF. (c) Ceiling analysis of the error sources.
The simulated population distributions shown in Fig. A6(a) are consistent with the experimental results shown in Fig. 2(c). In particular, vanilla SB’s result shows a nearly even distribution for the off-target states, while SPF’s one exhibits a distinct peak at $|N-1\rangle$. This can be explained as follows: although the SFP effectively corrects most transmon errors, the lengthy readout time for each transmon (approximately 1700 ns) involved in the correction process makes uncorrected qudit decoherence errors more prominent. In contrast to the photon-number insensitive transmon errors, which result in nearly uniform off-target populations, the effective error rates for qudits tend to increase with the photon number. Consequently, the off-target distribution is skewed towards $N$. In the simulation, the $|N-1\rangle$ peak is completely suppressed by PF, while it remains visible in the experimental data. This discrepancy arises because we do not account for the parity filtering efficiency in the simulation.
The predicted fidelities shown in Fig. A6(b) are higher than the experimental results presented in Fig. 2(d). This discrepancy is expected because our noise model is incomplete. In particular, control errors are not calibrated, and thus are not included in the noise model for simulation. The absence of control errors in the noise model is likely the reason why the trend of the predicted SB’s fidelities decreases at a slower rate than linear with increasing photon numbers. In contrast, the experimental fidelities show a faster-than-linear decrease as the photon number increases. In the simulation, the decoherence errors are approximately proportional to the sideband transition time, which decreases as $1/|\chi_{e}|\sim 1/\sqrt{N}$. However, as we move to higher photon numbers, we expect the control errors to become more pronounced as a result of the increasingly crowded energy levels. Therefore, it is likely that the control errors account for the faster-than-linear decrease in the experimental fidelities.
As we discussed earlier, SFP effectively corrects nearly all transmon decoherence errors, albeit at the cost of increasing qudit decoherence errors. This understanding is supported by the ceiling analysis illustrated in Figure A6(c). In our ceiling analysis, we toggle specific noise channels on and off. For instance, the red bar demonstrates the difference in fidelities when transmon pure dephasing is on compared to when it is off. This approach allows us to estimate the contributions of transmon pure dephasing and other noise sources to the overall infidelities. Comparing the SFP ceiling analysis on the right with the SB one on the left, the analysis clearly shows the effectiveness of SFP in correcting the transmon decoherence errors (red and green). Meanwhile, readout error (purple) and qudit depolarization (blue) become the limiting factors in the Fock state preparation protocol with SFP. This indicates that achieving higher fidelities and preparing higher Fock states will require further improvements in the ancilla transmon readout fidelity and the qudit $T_{1}$, in addition to the control errors mentioned above.
Appendix I Raman-assisted beamsplitter interaction
Here we derive the Alice-Bob beamsplitter interaction in the virtual Raman protocol. The Hamiltonian of the driven system is
$\hat{\mathcal{H}}$$\hbar$ $=\omega_{a}\hat{a}^{\dagger}\hat{a}+\omega_{b}\hat{b}^{\dagger}\hat{b}+\omega_{q0}\hat{q}^{\dagger}\hat{q}$
$-E_{J}\cos_{\mathrm{nl}}\left[\theta_{q}\left(\hat{q}+\frac{g_{a}}{\Delta_{a}}\hat{a}+\frac{g_{b}}{\Delta_{b}}\hat{b}\right)+\mathrm{h.c.}\right]$
$+\sum_{k=1,2}\epsilon_{k}\cos\left(\omega_{k}t\right)\left(\hat{q}+\hat{q}^{\dagger}\right),$ (I.1)
where $g_{a(b)}$ and $\Delta_{a(b)}$ stand for the coupling strength and detuning between Alice (Bob) and the transmon, and $\cos_{\mathrm{nl}}$ stands for 4-th and higher-order components of the cosine function. $\epsilon_{1(2)}$ and $\omega_{1(2)}$ are the amplitudes and frequencies of the two sideband drives acting on the transmon. Going into the displaced frame with respect to the sideband drives,
$\hat{\mathcal{H}}_{\mathrm{disp}}/\hbar=\omega_{a}\hat{a}^{\dagger}\hat{a}+\omega_{b}\hat{b}^{\dagger}\hat{b}+\omega_{q0}\hat{q}^{\dagger}\hat{q}$
$-E_{J}\cos_{\mathrm{nl}}\left[\theta_{q}\left(\hat{q}+\sum_{k=1,2}\xi_{k}e^{-i\omega_{k}t}+\frac{g_{a}}{\Delta_{a}}\hat{a}+\frac{g_{b}}{\Delta_{b}}\hat{b}\right)+\mathrm{h.c.}\right],$ (I.2)
where $\xi_{1,2}$ represents the drive-induced displacements. We further set the drive frequencies to be $\omega_{1(2)}=2\omega_{q0}+\alpha-\omega_{a(b)}-\Delta$ to match the VRBS resonance condition, where $\alpha$ is the transmon anharmonicity, and $\Delta$ is the virtual level detuning from $\lvert f00\rangle$. From going to the rotating frame and collecting the relevant slow-oscillating terms in $\cos_{\mathrm{nl}}$ expansion, we may approximate the virtual Raman (VR) Hamiltonian as
$\hat{\mathcal{H}}_{\mathrm{VR}}/\hbar\approx$ $\frac{\alpha}{2}\hat{q}^{\dagger}\hat{q}^{\dagger}\hat{q}\hat{q}+\chi_{a}\hat{q}^{\dagger}\hat{q}\hat{a}^{\dagger}\hat{a}+\chi_{b}\hat{q}^{\dagger}\hat{q}\hat{b}^{\dagger}\hat{b}$ (I.3)
$-\frac{E_{J}\theta_{q}^{4}}{2}\hat{q}^{\dagger 2}\left(\frac{\xi_{1}g_{a}}{\Delta_{a}}\hat{a}+\frac{\xi_{2}g_{b}}{\Delta_{b}}\hat{b}\right)e^{-i\Delta t}$
$-\frac{E_{J}\theta_{q}^{4}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{\Delta_{a}\Delta_{b}}\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.},$
where the third line represents the direct Alice-Bob beamsplitter interaction mediated by the 4th-order nonlinearity. By doing second-order perturbation theory to eliminate the virtual states, we find the total effective beamsplitter interaction between Alice and Bob in the $\{|10\rangle,|01\rangle\}$ subspace to be
$\hat{\mathcal{H}}_{\mathrm{VR}}^{\mathrm{BS}}/\hbar$ $\approx\left(\frac{E_{J}^{2}\theta_{q}^{8}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{2\Delta_{a}\Delta_{b}\Delta}-\frac{E_{J}\theta_{q}^{4}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{\Delta_{a}\Delta_{b}}\right)\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.}$ (I.4)
$\approx 2\left(\frac{\alpha}{\Delta}+1\right)\frac{\alpha\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{\Delta_{a}\Delta_{b}}\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.}.$ (I.5)
This clearly shows that, compared to the direct four-wave-mixing beamsplitter rate (represented by the 1 term), the VR scheme possesses a significant enhancement factor of $\alpha/\Delta$, making it over an order of magnitude faster than the former scheme in our experiment.
While the above derivation captures the essential interaction underlying the VR scheme, accurately modeling the system dynamics requires incorporating higher-order corrections induced by the drive or nonlinearity, which we demonstrate below. We apply a Schrieffer-Wolff transformation [Shavitt1980] with a generator of $\hat{\mathcal{S}}=\sum_{n=0,m=0}(\hat{\mathcal{S}}_{n,m}^{(a)}+\hat{\mathcal{S}}_{n,m}^{(b)}$), where
$\hat{\mathcal{S}}_{n,m}^{(a)}=$ $\frac{E_{J}\theta_{q}^{4}g_{a}\sqrt{6(n+1)}}{2\Delta_{a}[\Delta+\alpha+n\chi_{h}^{(a)}-(n+1)\chi_{e}^{(a)}+m\chi_{h}^{(b)}-m\chi_{e}^{(b)}]}|e,n+1,m\rangle\langle h,n,m|e^{i[\Delta+\alpha+n\chi_{h}^{(a)}-(n+1)\chi_{e}^{(a)}+m\chi_{h}^{(b)}-m\chi_{e}^{(b)}]t}$ (I.6)
$+\frac{E_{J}\theta_{q}^{4}g_{a}\sqrt{2(n+1)}}{2\Delta_{a}[\Delta+n\chi_{f}^{(a)}+m\chi_{f}^{(b)}]}|g,n+1,m\rangle\langle f,n,m|e^{i[\Delta+n\chi_{f}^{(a)}+m\chi_{f}^{(b)}]t}-\text{h.c.},$
$\hat{\mathcal{S}}_{n,m}^{(b)}=$ $\frac{E_{J}\theta_{q}^{4}g_{b}\sqrt{6(m+1)}}{2\Delta_{b}[\Delta+\alpha+n\chi_{h}^{(a)}-n\chi_{e}^{(a)}+m\chi_{h}^{(b)}-(m+1)\chi_{e}^{(b)}]}|e,n,m+1\rangle\langle h,n+1,m|e^{i[\Delta+\alpha+n\chi_{h}^{(a)}-n\chi_{e}^{(a)}+m\chi_{h}^{(b)}-(m+1)\chi_{e}^{(b)}]t}$ (I.7)
$+\frac{E_{J}\theta_{q}^{4}g_{b}\sqrt{2(m+1)}}{2\Delta_{b}[\Delta+n\chi_{f}^{(a)}+m\chi_{f}^{(b)}]}|g,n,m+1\rangle\langle f,n+1,m|e^{i[\Delta+n\chi_{f}^{(a)}+m\chi_{f}^{(b)}]t}-\text{h.c.}.$
Here, $\chi^{(a,b)}_{i}$ denotes the dispersive shift of the Alice (a) or Bob (b) cavity mode conditioned on the ancilla being in state $|i\rangle$. Applying the Baker–Campbell–Hausdorff formula, the second line of Eq. (I.3) transforms into
$\frac{1}{2}\left[\hat{\mathcal{S}},-\frac{E_{J}\theta_{q}^{4}}{2}\hat{q}^{\dagger 2}\left(\frac{\xi_{1}g_{a}}{\Delta_{a}}\hat{a}+\frac{\xi_{2}g_{b}}{\Delta_{b}}\hat{b}\right)e^{-i\Delta t}+\text{h.c.}\right]=$
$\sum_{n,m}\beta_{a}\beta_{b}\sqrt{n{+}1}\sqrt{m{+}1}|g\rangle\langle g|\otimes|n,m{+}1\rangle\langle n{+}1,m|\left(\frac{2}{\Delta+n\chi_{f}^{(a)}+m\chi_{f}^{(b)}}\right)$
$-\beta_{a}\beta_{b}\sqrt{n{+}1}\sqrt{m{+}1}|f\rangle\langle f|\otimes|n,m{+}1\rangle\langle n{+}1,m|e^{i(\chi_{f}^{(a)}-\chi_{f}^{(b)})t}\left[\frac{1}{\Delta+n\chi_{f}^{(a)}+(m{+}1)\chi_{f}^{(b)}}+\frac{1}{\Delta+(n{+}1)\chi_{f}^{(a)}+m\chi_{f}^{(b)}}\right]$
$+3\beta_{a}\beta_{b}\sqrt{n{+}1}\sqrt{m{+}1}|e\rangle\langle e|\otimes|n,m{+}1\rangle\langle n{+}1,m|e^{i(\chi_{e}^{(a)}-\chi_{e}^{(b)})t}$
$\times\left[\frac{1}{\Delta+\alpha+n\chi_{h}^{(a)}-(n{+}1)\chi_{e}^{(a)}+m\chi_{h}^{(b)}-m\chi_{e}^{(b)}}+\frac{1}{\Delta+\alpha+n\chi_{h}^{(a)}-n\chi_{e}^{(a)}+m\chi_{h}^{(b)}-(m{+}1)\chi_{e}^{(b)}}\right]$
$-3\beta_{a}\beta_{b}\sqrt{n{+}1}\sqrt{m{+}1}|h\rangle\langle h|\otimes|n,m{+}1\rangle\langle n{+}1,m|e^{i(\chi_{h}^{(a)}-\chi_{h}^{(b)})t}$
$\times\left[\frac{1}{\Delta+\alpha+(n{+}1)\chi_{h}^{(a)}-(n{+}1)\chi_{e}^{(a)}+(m{+}1)\chi_{h}^{(b)}-(m{+}1)\chi_{e}^{(b)}}+\frac{1}{\Delta+\alpha+(n{+}1)\chi_{h}^{(a)}-(n{+}1)\chi_{e}^{(a)}+m\chi_{h}^{(b)}-(m{+}1)\chi_{e}^{(b)}}\right]$
$+\text{h.c.},$
where we define the quantities $\beta_{a,b}=\frac{E_{J}\theta_{q}^{4}\xi_{1,2}g_{a,b}}{2\Delta_{a,b}}$. From the equation above, we note that the beasmplitter rate is highly dependent on the transmon state, with detunings in the energy denominator differing by the anharmonicity for transmon $\lvert g\rangle$ and $\lvert e\rangle$ states. In general, the dispersive shifts are different for the two cavities. When the transmon is in the ground state, the resonant interaction is
$\sum_{n,m}\frac{E_{J}^{2}\theta_{q}^{8}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{2\Delta_{a}\Delta_{b}}\left(\frac{|g\rangle\langle g|}{\Delta+n\chi_{f}^{(a)}+m\chi_{f}^{(b)}}\right)\otimes\sqrt{(n+1)(m+1)}(|n,m+1\rangle\langle n+1,m|+|n+1,m\rangle\langle n,m+1|).$ (I.8)
This expression can be further simplied in the regime where $\chi_{f}^{(a,b)}\ll\Delta$:
$\frac{E_{J}^{2}\theta_{q}^{8}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{2\Delta\Delta_{a}\Delta_{b}}|g\rangle\langle g|\otimes(\hat{a}\hat{b}^{\dagger}+\hat{a}^{\dagger}\hat{b})\left[1-\frac{1}{\Delta}(\hat{n}_{a}\chi_{f}^{(a)}+\hat{n}_{b}\chi_{f}^{(b)})\right].$ (I.9)
Intuitively, the beamsplitter rate is dependent on the cavity photon number because the drive detuning to the virtual states (VRBS detuning) for higher-photon-number sectors is altered by their dispersive shifts. This nonlinear correction to the beamsplitter gate has applications in universal qudit control, as discussed in Appendix N.
Figure A12 shows a representative example of the VRBS operation in matrix form, where the effects on higher Fock states are obtained by numerically solving the Schrödinger equation (left panel). The block-diagonal structure of the evolution operator can be understood analytically from Eq. (I.3), whose corresponding unitary evolution is shown in the right panel. The analytic result shows good agreement with the exact numerical simulation, with small deviations likely resulting from higher-order processes that are neglected in the derivation.
Appendix J Numerical simulation of VR entanglement
Figure A7: Numerical simulations of Raman-assisted beamsplitter interactions. (a) Simulated $|f0\rangle\leftrightarrow|g1\rangle$ sideband dynamics between Alice and the ancilla as a function of drive detuning with respect to the sideband resonance. (b) Line cuts taken along the resonant sideband condition indicated by the red dashed line in (a). (c) Simulated Raman-assisted population exchange between Alice and Bob, mediated by the ancilla and driven by dual sideband tones. (d) Line cuts of the cavity population dynamics at representative detunings. The dashed curve shows the transmon ground state population at a detuning of 1.3-1.3 MHz, which aligns with the modulation observed in the corresponding cavity population.
Here, we provide details on the numerical simulations of the Raman-assisted beamsplitter dynamics. The Hamiltonian in the bare basis is
[…]
$+\sum_{k=1,2}\epsilon_{k}\cos\left(\omega_{k}t\right)\left(\hat{q}+\hat{q}^{\dagger}\right).$
The circuit parameters are chosen such that the static energy spectrum of the Hamiltonian reproduces the experimental values listed in Table A1.
Next, we calibrate the sideband drives for Alice and Bob individually. For a fixed sideband amplitude (e.g., $\epsilon_{1}$), we sweep the corresponding drive frequency $\omega_{1}$ to extract the resonant $|f0\rangle$$|g1\rangle$ sideband oscillation period between Alice and the ancilla. This procedure is repeated until the simulated resonant oscillation period matches the experimentally observed value, thereby determining both the sideband drive amplitude $\epsilon_{1}$ and frequency $\omega_{1}$. As an example, Fig. A7(a) shows the simulated sideband dynamics for Alice and ancilla. The red lines indicate the resonant conditions, and the corresponding frequency line cuts are shown in Fig. A7(b).
With these calibrated parameters, we simulate the system evolution under simultaneous sideband drives, with identical detunings from their respective resonances. Figure A7(c) displays the resulting population exchange between Alice and Bob, conditioned on the ancilla being in the ground state. For illustration, Fig. A7(d) shows representative line cuts at several detunings. At large detunings, the population exhibits sinusoidal oscillations. As the detuning decreases, the oscillation frequency increases and develops a high-frequency modulation. This modulation originates from population in the transmon’s $|f\rangle$ state, due to the breakdown of the adiabatic condition at small sideband detuning. For instance, the transmon ground state population at a detuning of 1.3-1.3 MHz, shown as a dashed line, aligns with the modulation observed in the corresponding beamsplitter dynamics. Figure 3(b) plots the extracted beamsplitter rates as a function of detuning (red line), showing excellent agreement with experimental data (blue dots).
While the above simulations capture the coherent dynamics, we also perform open-system simulations to study decoherence and its effect on gate fidelity. These are based on solving the Lindblad master equation with the following bare-basis Lindbladian:
$\hat{\mathcal{L}}[\rho]/\hbar=$ $-i[\hat{H},\rho]+(T_{1}^{ge})^{-1}\hat{\mathcal{D}}[\hat{q}]\rho$ (J.2)
$+\bar{n}^{\text{q}}_{\text{th}}(T_{1}^{ge})^{-1}\hat{\mathcal{D}}[\hat{q}^{\dagger}]\rho+2(T_{\phi}^{ge})^{-1}\hat{\mathcal{D}}[\hat{q}^{\dagger}\hat{q}]\rho$
$+(T_{1}^{\text{A}})^{-1}\hat{\mathcal{D}}[\hat{a}]\rho+(T_{1}^{\text{B}})^{-1}\hat{\mathcal{D}}[\hat{b}]\rho.$
[…]
Figure A8: VRBS drive calibration (a) Bob’s Fock $|1\rangle$ state population $P_{1}^{B}$ as a function of Bob sideband drive frequency with fixed duration, swept around the transition frequency $\omega_{|f00\rangle\leftrightarrow|g01\rangle}+\Delta$. The peak frequency corresponds to the difference in Stark shifts between the Alice and Bob modes. (b) Accumulated phase over multiple VRBS swap operations, measured by sweeping the phase of the final $U_{\text{BS}}(\phi)$ and identifying the phase correction that maximizes target state fidelity.
[…]
Appendix K VRBS drives calibration
In this section, we discuss the calibration procedure for the VRBS drives. Accurate VRBS operation critically depends on precisely satisfying the frequency matching condition. Due to the drive-induced Stark shifts, the effective detuning of the VRBS drives differs from the detuning between Alice and Bob in the absence of driving. To account for this effect, we first prepare the Alice mode in the state $|1\rangle$. Subsequently, we simultaneously apply the $|f00\rangle\leftrightarrow|g10\rangle$ and $|f00\rangle\leftrightarrow|g01\rangle$ sideband drives with same detuning at appropriate drive amplitudes. We keep the Alice sideband drive frequency fixed while sweeping the Bob sideband drive frequency. The precise resonance condition is then detected as the peak in the Bob $|1\rangle$ state population measurement (see Fig. A8(a)).
The small but finite difference in Alice’s and Bob’s drive-induced Stark shifts results in a relative phase rotation between the Alice and Bob modes, accumulated over idling between concatenated VRBS pulses or ramping during the pulses [LuSchoelkopf2023]. For the error-detected VRBS swap operation, the transmon measurement after each swap can also contribute to a differential phase rotation of Alice and Bob due to their different cross-Kerr values with the readout mode (see Table A1). Without proper calibration and compensation, these accumulated phases will translate to a coherent error that reduces the fidelity of the $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$ state swap operation. In our experiment, we measure the phase rotation accumulated over a number of VRBS swap operations, by initializing the system in the $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$ state, and varying the phase $\phi$ of the final VRBS beamsplitter operation that maps the entangled state back to the single-photon Fock state in Alice. The accumulated phase is identified by the angle that maximizes the target state fidelity. We observe a linear increase of the accumulated phase with respect to the number of swap operations, where we extract the phase shift per gate to be $0.0046$ rad, see Fig. A8(b). This phase $\phi$ is then applied to the beamsplitter operation $U_{\text{BS}}(\phi)$ for the error-detected swap operation of the state $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$.
Appendix L Transmon heating detection after each swap operation
Transmon heating events during swap operation prevent coherent oscillation between Alice and Bob, thereby causing dephasing. To detect and characterize these heating events, we measure the transmon state immediately after each swap operation. We perform measurements up to 100 consecutive swap operations, post-selecting out any instances where heating is detected. We fit the transmon heating detection probability as a function of number of swaps, $P(n)=1-(1-P_{\uparrow})^{n}$, to estimate the transmon heating rate per swap operation, $P_{\uparrow}$. Figure A9 illustrates the heating detection rate observed over these 100 swap operations. From these measurements, we estimate a heating probability of approximately $P_{\uparrow}\approx 1.167\pm 0.003\%$ per swap operation. This heating (and decay) of the driven ancilla results in stochastic hopping of the ancilla population between the ground and excited state, further leading to the dephasing of the VRBS process due to the large dispersion of the effective VRBS rate over different ancilla states (see Appendix I). Therefore, by adding mid-circuit transmon measurements, we are able to detect the dominant dephasing error in VRBS, and demonstrate an improvement in the VRBS fidelity by post-selecting on no heating events, see main text Fig. 3.
Figure A9: Transmon heating during VRBS drive (a) Transmon population during continuous VRBS drive. The transmon heats up, reaching a thermal population of $\sim 4\%$ after a 1 ms VRBS drive. (b) The probability of the transmon heating event up to 100 swap gate operations. We estimate the transmon heating probability as $1.167\pm 0.003$ % per swap gate operation.
Figure A10: Cavity $T_{1}$ as a function of transmon measurement counts. (a) The quantum circuit showing the experimental protocol for measuring the cavity $T_{1}$ in the presence of multiple transmon readouts during the wait period. (b) The relaxation times are not affected by transmon readouts.
[…]
Appendix M Mapping dual-rail cavity states onto transmon states
Figure A11: Mapping of cavity states to the transmon. The transmon is first reset to the $|g\rangle$ state by measurement and following conditional $\pi$ pulse(s). Then a combination of sideband and transmon pulses are applied to perform the mapping: $|00\rangle\rightarrow|g\rangle,\ |10\rangle\rightarrow|e\rangle$ and $|01\rangle\rightarrow|f\rangle$.
Figure A12: Matrix elements of the VRBS operation. The matrix takes a block-diagonal form, reflecting the structure of the VRBS gate when acting on higher Fock states. The left panel shows the exact numerical result obtained by solving the time-dependent Schrödinger equation, while the right panel displays the corresponding analytical unitary evolution derived from Eq. (I.3). The two results show good agreement. The simulation is performed at a sideband detuning of 5.5-5.5 MHz. Only the probability amplitudes are shown for brevity.
To measure the populations of single-photon states in the Alice and Bob modes using single-shot readout, we map the cavity states onto the transmon states. The circuit diagram is illustrated in Fig. A11. After reseting the transmon to the ground state, we apply a sideband transition pulse $|g10\rangle\leftrightarrow|f00\rangle$ (corresponding to Alice mode). Subsequently, we apply a photon-number-unselective $\pi_{ef}$ pulse with a short duration ($\Delta t\ll 2\pi/\chi_{e}^{a},\,2\pi/\chi_{e}^{b}$). Finally, we apply a second sideband transition pulse $|g01\rangle\leftrightarrow|f00\rangle$ (corresponding to Bob mode). Through this sequence, the population in the $|10\rangle$ state maps to the transmon $|e\rangle$ state, the $|01\rangle$ state population maps to $|f\rangle$, and the $|00\rangle$ state maps to $|g\rangle$. This method enables single-shot measurement of $|00\rangle,\ |01\rangle,\ |10\rangle$ states within a duration significantly shorter ($\ll 1/\chi_{e}$) compared to parity measurement or PNRS, which typically have durations on the order of $1/\chi_{e}$. The resulting confusion matrix $M_{\rm map}$ for this mapping is
$M_{\rm map}=\begin{bmatrix}0.9959&0.0491&0.0100\\ 0.0018&0.9467&0.0172\\ 0.0022&0.0042&0.9728\end{bmatrix}$, (M.1)
where the columns (rows) correspond to the prepared (measured) states of $|00\rangle,|10\rangle$ and $|01\rangle$. The observed asymmetry between Alice and Bob arises from the order of the mapping procedure. We use this confusion matrix to correct the cavity state readout errors for the beamsplitter operations.
Appendix N Qudit gates
Figure A13: Simplified cryogenic wiring diagram.
Universal control of a two-qudit system requires the ability to perform arbitrary single-qudit rotations along with at least one entangling operation [Braunstein2005CV]. In our multimode architecture, single-qudit rotations $R_{d}(\vec{\theta})$ can be implemented using either sideband gates [Liu2021qudit] or echoed conditional displacement gates [You2024crosstalk]. The VRBS interaction discussed in the main text provides a mechanism for generating entanglement between higher Fock states as well. This operation takes a block-diagonal form, where each basis state $|m,n\rangle$ is rotated within a photon-number-conserving subspace—i.e., among states $|m^{\prime},n^{\prime}\rangle$ satisfying $m^{\prime}+n^{\prime}=m+n$. Together with single-qudit rotations, this operation enable universal control within a finite-dimensional Hilbert space constructed from $d$-dimensional qudits. For example, universal operations in the two-qudit Hilbert space can be performed by:
$\mathcal{U}=\textsc{R}_{d}(\vec{\theta}_{0})\otimes\textsc{R}_{d}(\vec{\phi}_{0})\prod_{i=1}^{N_{\rm block}}\left[\textsc{VRBS}\;\textsc{R}_{d}(\vec{\theta}_{i})\otimes\textsc{R}_{d}(\vec{\phi}_{i})\right],$ (N.1)
where $N_{\rm block}$ is the number of blocks of those gates within the square brackets.
With the results of Fig. A12, it is possible to investigate the properties of VRBS as a two-qutrit entangling gate, as a concrete example. For qutrits, it has a Schmidt rank of 9. The entangling power, $e_{p}(U)$ can be defined as the average of the linear entropy over product states $|\Psi\rangle=|\psi\rangle_{1}\otimes|\psi_{2}\rangle$ following [wang2003entangling]:
$E(|\Psi\rangle)=1-\text{Tr}_{1}\rho_{1}^{2},$ (N.2)
where $\rho_{1}=\text{Tr}_{2}(U|\Psi\rangle\langle\Psi|U^{\dagger})$ is the reduced density matrix, and $U$ is the entangling gate being investigated. We estimate $e_{p}(\rm VRBS)=0.379(2)$ by sampling $10^{5}$ $|\Psi\rangle$, similar to the standard qutrit CSUM gate (${\rm CSUM_{3}}|m,n\rangle=|m,m+n\ {\rm mod}\ 3\rangle$) which is known analytically to have $e_{p}(\text{CSUM}_{3})=3/8$ [wang2003entangling]. This implies that VRBS is an efficient qutrit entangling gate. Further, by numerical optimization of the $SU(3)$ angles in Eq. (N.1), approximate synthesis of the CSUM3 gate from $2\leq N\leq 6$ blocks were identified with fidelities
$\mathcal{F}_{\text{CSUM}_{3}}^{N_{\rm blocks}}=\{51\%,88\%,92\%,99.1\%,99.999974\%\}.$ (N.3)
Demonstration of universality in our 2-cell system is the topic of future work, with emphasis placed on qudit CSUM gates.
Appendix O Experimental setup
Figure A13 illustrates the cryogenic wiring diagram of our measurement setup. The microwave components and design choices are carefully optimized to minimize thermal excitation of both the cavity modes and the transmon. The double-cell cavity is mounted vertically within a custom-built support structure—or cage—comprising straight and semicircular clamps made from oxygen-free high thermal conductivity (OFHC) copper, ensuring efficient thermal anchoring and heat extraction.
To suppress infrared (IR) radiation, the cage is wrapped in EccosorbTM foam, which absorbs stray IR photons that could otherwise be absorbed by the cavity or surrounding microwave components. Two commercial Eccosorb filters, thermalized via OFHC copper braids, are installed on the drive lines to the cavity and transmon, providing attenuation of high-energy IR photons coming down the coaxial cables. A High-Energy-Radiation-Drain (HERD-2) filter is placed on the readout line to block infrared radiation while minimizing signal attenuation, ensuring high-fidelity readout. All IR filters are mounted inside the Eccosorb-wrapped region for protection against ambient IR photons.
A copper can is mounted on the dilution refrigerator’s base plate to enclose and shield all components thermalized to the base temperature at around 8 mK. To further reduce susceptibility to magnetic noise, a $\mu$-metal shield is placed around the vacuum chamber, protecting the system from external magnetic fields.
For control and measurement, we use the OPX+ and Octave RF hardware from Quantum Machines. The transmon is read out in reflection via its on-chip readout resonator. A Josephson Parametric Amplifier (JPA) is used as the first-stage phase-insensitive amplifier in the readout chain. The JPA’s charge pump (8.375 GHz) and DC bias are supplied by a Rhode&Schwarz SMA100B signal generator and a Yokogawa GS200 DC voltage/current source, respectively.